Oxygen at 5 bar and 320 K is supplied to an insulated tank of volume 5

m3

. The tank initially contains oxygen at 1 bar and 300 K. Determine:

i) the final temperature in the tank when its pressure reaches 5 bar. 

ii) the mass that has entered the tank.


1
Expert's answer
2021-06-16T14:23:11-0400

Gives

"P_1=100kpa"

T1=300k

"V_1=\frac{P_2V_2 \times T_1}{T_2\times P_1}"

Put Value

V1=23.43 m3

"P_2=500kpa"

"T_2=320K"

"V_2=5m^3"

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"

"T_2=\frac{T_1}{P_1V_1}\times P_2V_2"

Put value

"T_2=\frac{300 \times500Kpa\times5}{100Kpa\times23.43}="320K

Part(b)

We know that

"m=\frac{P_1V_1}{RT_1}"

"m=\frac{100 \times23.43}{4.12\times400}=1.42kg"